### v2

# A semicircle law for the normalized Laplacian of sparse random graphs

Abstract.
We study the limiting spectral distribution of the normalized Laplacian $\mathcal{L}$ of an Erdős-Rényi graph $G(n,p)$. To account for the presence of isolated vertices in the sparse regime, we define $\mathcal{L}$ using the Moore-Penrose pseudoinverse of the degree matrix. Under this convention, we show that the empirical spectral distribution of a suitably normalized $\mathcal{L}$ converges weakly in probability to the semicircle law whenever $np\to\infty$, thereby providing a rigorous justification of a prediction made in (Akara-pipattana and Evnin, 2023). Moreover, if $np>\log n+\omega(1)$, so that $G(n,p)$ has no isolated vertices with high probability, the same conclusion holds for the standard definition of $\mathcal{L}$. We further strengthen this result to almost sure convergence when $np=\Omega(\log n)$. Finally, we extend our approach to the Chung-Lu random graph model, where we establish a semicircle law for $\mathcal{L}$ itself, improving upon (Chung, Lu, and Vu 2003), which obtained the semicircle law only for a proxy matrix.
1. Introduction
Spectral questions for random graphs lie at the interface of random matrix theory and random graph theory, and play a central role in a wide range of applications, including high-dimensional statistics [le2018concentration] and theoretical computer science [hoory2006expander]. Given a graph, the eigenvalues and eigenvectors of its associated matrices control fundamental structural and algorithmic phenomena [chung1997spectral] such as expansion and connectivity, mixing and hitting times of random walks, and the performance of spectral algorithms for clustering and community detection. A canonical and extensively studied model is the Erdős–Rényi graph $G(n,p)$, whose adjacency matrix provides a prototypical example of a sparse random matrix. Over the past decades, global and local eigenvalue statistics, spectral norm bounds, and eigenvector properties for adjacency matrices have been developed across a broad range of sparsity regimes; see, e.g., [erdHos2013spectral, tran2013sparse, benaych2019largest, alt2021extremal] and references therein.
In spectral graph theory [chung1997spectral], two particularly important operators are the (combinatorial) Laplacian $\mathbf{L}\;=\;\mathbf{D}-\mathbf{A}$, and the normalized Laplacian
$\mathbf{\mathcal{L}}\;=\;\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2},$ (1)
where $\mathbf{A}$ is the adjacency matrix and $\mathbf{D}$ is the diagonal degree matrix. These matrices are ubiquitous in both theory and algorithms: $\mathbf{L}$ is the basic object for cuts, connectivity, and electrical network interpretations [lyons2017probability], while $\mathbf{\mathcal{L}}$ (and the closely related random-walk matrix $\mathbf{D}^{-1}\mathbf{A}$) governs random-walk behavior and underpins spectral clustering [meilua2001random]. Compared with adjacency matrices, however, the spectral theory of Laplacian-type matrices for random graphs is less complete, especially in regimes where degree fluctuations are significant. The normalization in (1) introduces a random and dependent reweighting through $\mathbf{D}^{-1/2}$, creating challenges that do not arise in the spectral analysis of $\mathbf{A}$.
Laplacian random matrices
For the Laplacian $\mathbf{L}$ of an Erdős–Rényi graph, the empirical spectral distribution (after appropriate centering and scaling) has been shown to converge to the free convolution of a Gaussian distribution with the semicircle law as long as $np\to\infty$; see, e.g., [jiang2012empirical, ding2010spectral, huang2020spectral, chakrabarty2021spectra, chatterjee2022spectral, campbell2024spectrum]. Beyond global limits, the behavior of extreme eigenvalues (or singular values) has also attracted considerable attention, with recent works establishing sharp asymptotics and fluctuation results in various regimes [campbell2025extreme, ducatez2025spectral, christoffersen2024eigenvalue, subhra2025large]. Related deformed Laplacian matrices, including the Bethe-Hessian, have been studied in the context of community detection in stochastic block models [saade2014spectral, dall2021unified, stephan2024community]. For directed sparse random graphs, limiting spectral laws and related phenomena have also been investigated; see [bordenave2014spectrum].
The normalized Laplacian
A useful perspective on $\mathbf{\mathcal{L}}$ comes from random walks: writing
$\mathbf{\mathcal{M}}\;:=\;\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2},\qquad\mathbf{\mathcal{L}}\;=\;\mathbf{I}-\mathbf{\mathcal{M}},$
the matrix $\mathbf{\mathcal{M}}$ is a symmetric (reversible) Markov kernel obtained by symmetrizing the random-walk matrix $\mathbf{D}^{-1}\mathbf{A}$. Since the spectra of $\mathcal{L}$ and $\mathcal{M}$ have a deterministic relation, we will focus on the spectrum of $\mathcal{M}$ below.
In dense regimes, since $\mathbf{D}$ is concentrated around $(np)\mathbf{I}$, the semicircular law for $\mathcal{M}$ is expected. When $np=\omega(\log n)$, Jiang [jiang2012empirical] showed that the empirical spectral distribution of $\sqrt{np},\mathcal{M}$ converges almost surely to the semicircle law. In contrast, in the bounded expected degree regime, for a regularized version of $\mathcal{M}$, Chi [chi2016random] identified a non-degenerate symmetric limiting distribution characterized by return probabilities on Galton–Watson trees for $G(n,c/n)$.
In this paper, we bridge the gap between the constant expected-degree regime $(np=c)$ and the supercritical regime $(np=\omega(\log n))$. We show that, as long as $np\to\infty$, the semicircle law for $\mathcal{M}$ continues to hold, provided one adopts a convention for $\mathbf{D}^{-1/2}$ that remains well defined in the presence of isolated vertices (equivalently, by projecting away zero-degree coordinates). In view of the result of Chi [chi2016random], the condition $np\to\infty$ is both necessary and sufficient for the semicircle law to hold. The semicircle law for $\mathcal{M}$ when $np\to\infty$ was also conjectured by Akara-pipattana and Evnin in [akara2023random, Section 5.2] based on predictions from the Fyodorov–Mirlin method, and we provide a rigorous justification.
Complementing asymptotic results, a substantial line of work establishes non-asymptotic extreme eigenvalue bounds of $\mathbf{\mathcal{L}}$, motivated in part by stability guarantees for spectral clustering [le2017concentration, deng2021strong, chen2025concentration]; see also [hoffman2021spectral] for a sharp spectral-gap estimate restricted to the giant component of $G(n,p)$. Beyond random graph models, the limiting spectrum of the random walk matrix $\mathbf{D}^{-1}\mathbf{A}$ has also been considered for more general Hermitian [bordenav2011spectrum, bordenave2010spectrum] or non-Hermitian random matrix $\mathbf{A}$ [bordenave2012circular].
The Chung-Lu model
Going beyond homogeneous Erdős–Rényi graphs, inhomogeneous random graphs $G(n,p_{ij})$, where edges $(i,j)$ are present independently with probabilities $p_{ij}$, provide a flexible modeling framework that captures heterogeneous expected degrees and latent structure. A prominent special case is the rank-one (Chung-Lu) model [chung2002average, chung2006complex], which prescribes expected degrees via a factorized probability matrix. See Section 2.2 for a detailed description of the model. In this setting, Chung, Lu, and Vu [chung2003spectra] established a semicircle law for a matrix $\overline{\mathcal{M}}=\mathbb{E}[\mathbf{D}]^{-1/2}\mathbf{A}\mathbb{E}[\mathbf{D}]^{-1/2}$ approximating the normalized Laplacian, and further spectral estimates of $\mathcal{L}$ were developed in [lu2013spectra]. Local spectral statistics for its adjacency matrix were investigated in [adlam2015spectral].
We extend the semicircle law of the normalized Laplacian for $G(n,p)$ to the Chung–Lu model. In particular, we show that after normalization, a semicircle law holds for $\mathcal{M}$ itself, rather than the proxy matrix $\overline{\mathcal{M}}$ studied in [chung2003spectra].
Beyond the rank-one setting, for general inhomogeneous random graphs, the limiting spectral distribution of $\mathcal{M}$ need not be semicircular, even in dense regimes. If one further assumes that the variance profile of $\overline{\mathcal{M}}$ admits a graphon limit [zhu2020graphon], the comparison argument developed in this paper can still be applied to characterize the corresponding limiting spectral distribution of $\mathcal{L}$.
Notation
Let $(\mathbb{R},\mathcal{B})$ denote the measurable space, where $\mathcal{B}$ is the Borel $\sigma$-algebra on $\mathbb{R}$. The indicator function is denoted by $\mathbf{1}(\cdot)$, and $\delta_{x}$ stands for the Dirac measure concentrated at $x\in\mathbb{R}^{d}$. The constant $c$ may change from line to line. We also use the notation $n]:=\{1,2,\dots,n\}$. For two sequences of nonnegative real numbers $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$, and $a_{n}=O(b_{n})$ if $a_{n}\leq Cb_{n}$ for some $C>0$ and all large $n$. We write $a_{n}\gg b_{n}$ if $a_{n}/b_{n}\to\infty$.
We denote by $\mu_{\mathrm{sc}}$ the semicircle distribution supported on [2,2][-2,2], with density function
$d\mu_{\mathrm{sc}}=\frac{1}{2\pi}\sqrt{4-x^{2}}\mathbf{1}\{|x|\leq 2\}dx.$ (2)
Organization of the paper
The remainder of this paper is organized as follows. We state our main results for the normalized Laplacian of $G(n,p)$ and the Chung-Lu model in Section 2. Section 3 includes some basic facts about random matrix theory. We prove Theorems 1 and 2 in Section 4 and the proof of Theorem 3 is given in Section 5.
2. Main Results
2.1. The Erdős-Rényi graph $G(n,p_{n})$
Let $\mathbf{A}_{n}=(a_{ij}^{(n)})$ be the adjacency matrix of an Erdős-Rényi graph $G(n,p_{n})$ such that $a_{ii}^{(n)}=0,i\in[n]$. The normalized Laplacian matrix is defined by
$\mathbf{\mathcal{L}}_{n}=\mathbf{I}_{n}-\mathbf{D}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{D}_{n}^{-1/2},$ (3)
where $\mathbf{D}_{n}=\operatorname{diag}(\sum_{j\neq i}^{n}a_{ij}^{(n)})_{1\leq i\leq n}$ is a diagonal matrix. If a graph has isolated vertices, $\mathbf{D}_{n}$ is not invertible. Still, we can extend the definition of $\mathbf{D}_{n}^{-1}$ using the Moore-Penrose pseudoinverse [van1996matrix] by defining $\mathbf{D}_{n}^{-1}(i,i)=0$ if $\mathbf{D}_{n}(i,i)=0$. Under this convention, we show that the semicircle law holds when $np_{n}(1-p_{n})\to\infty$.
Theorem 1 (Semicircle law for $\mathcal{L}_{n}$).
Assume $\mathbf{D}_{n}^{-1}$ is defined using the pseudo-inverse convention. Suppose that there exists a fixed constant $\varepsilon_{0}>0$ such that $\sup_{n}p_{n}\leq 1-\varepsilon_{0}$ and
$np_{n}\to\infty\quad\text{as }n\to\infty.$ (4)
Then the empirical spectral distribution of $\sqrt{\frac{np_{n}}{1-p_{n}}}(\mathbf{I}_{n}-\mathcal{L}_{n})$ converges weakly to the semicircle law $\mu_{\mathrm{sc}}$ defined in (2), in probability as $n\to\infty$.
Remark 1 (Optimality).
When $np_{n}=c$, [chi2016random, Theorem 5] showed that the limiting spectral distribution of $\mathbf{I}_{n}-\mathcal{L}_{n}$ is not semicircular. Consequently, the condition $np_{n}\to\infty$ is optimal for the semicircle law to hold, improving upon the stronger assumption $np_{n}=\omega(\log n)$ required in [jiang2012empirical]. We note that Theorem 1 was also conjectured in [akara2023random, Section 5] based on predictions from the Fyodorov–Mirlin method.
Since $\mathbf{D}_{n}$ is invertible if and only if $G(n,p_{n})$ has no isolated vertices, which occurs with high probability whenever $np_{n}>\log n+\omega(1)$ (see, e.g., [bollobas2011random, Theorem 7.3]), we obtain the following corollary of Theorem 1:
Corollary 1.
Assume that $\sup_{n}p_{n}<1-\varepsilon_{0}$ and there exists a sequence $c_{n}\to\infty$ such that
$np_{n}\geq\log n+c_{n}.$
Then, with probability tending to one, the degree matrix $\mathbf{D}_{n}$ is invertible. Moreover, the empirical spectral distribution of $\sqrt{\frac{np_{n}}{1-p_{n}}}(\mathbf{I}_{n}-\mathcal{L}_{n})$ converges weakly in probability to the semicircle law as $n\to\infty$.
By further assuming that $np_{n}=\Omega(\log n)$, we can strengthen convergence in probability to almost sure convergence, as stated below:
Theorem 2.
Suppose that $\sup_{n}p_{n}\leq 1-\varepsilon_{0}$, and $np_{n}\geq C\log n$ for some absolute constant $C$. Then, almost surely, the empirical spectral distribution of $\sqrt{\frac{np_{n}}{1-p_{n}}}(\mathbf{I}_{n}-\mathcal{L}_{n})$ converges weakly to the semicircle law.
A careful examination of the proof indicates that Theorem 2 holds for $C>36$, although no attempt is made to optimize the numerical constant.
2.2. The Chung–Lu Model
The Chung–Lu model [chung2002average, chung2006complex] is a widely used random graph model for generating random graphs with a prescribed expected degree sequence. It can be viewed as a natural generalization of the Erdős–Rényi random graph. Let $n\in\mathbb{N}$ and let
$\mathbf{w}_{n}=(w_{n,1},w_{n,2},\dots,w_{n,n})$
be a sequence of positive weights. Define
$\phi_{n}:=\frac{1}{\sum_{i=1}^{n}w_{n,i}}.$
The Chung–Lu random graph $G(n,\mathbf{w}_{n})$ is an undirected random graph on the vertex set $n]:=\{1,2,\dots,n\}$, in which edges between distinct vertices $i\neq j$ are placed independently with probabilities
$\mathbb{P}\!\left(a_{ij}^{(n)}=1\right)=p_{n,ij}:=w_{n,i}w_{n,j}\phi_{n},\qquad\mathbb{P}\!\left(a_{ij}^{(n)}=0\right)=1-p_{n,ij}.$
Here $a_{ij}^{(n)}$ denotes the adjacency indicator between vertices $i$ and $j$.
Let $\mathbf{A}_{n}=(a_{ij}^{(n)})_{1\leq i,j\leq n}$ denote the adjacency matrix of $G(n,\mathbf{w}_{n})$, and let $\mathbf{D}_{n}=\operatorname{diag}(d_{n,1},\dots,d_{n,n})$ be the corresponding degree matrix. We further denote by $\mathbf{W}_{n}$ the diagonal matrix with diagonal entries $w_{n,i}$, representing the expected degrees of the vertices. A fundamental property of the Chung-Lu model is that the expected degree of vertex $i$ satisfies
$\mathbb{E}[d_{n,i}]=\sum_{j=1}^{n}p_{n,ij}\approx w_{n,i},$
provided that
$\max_{1\leq i\leq n}w_{n,i}^{2}\;<\;\sum_{j=1}^{n}w_{n,j},$
which ensures that $p_{n,ij}\leq 1$ for all $i,j$. Consequently, the weight $w_{n,i}$ can be interpreted as the target expected degree of vertex $i$. As a special case, if $w_{n,i}\equiv np_{n}$ for all $i$, then $p_{n,ij}=p_{n}$, and the Chung-Lu model reduces to the classical Erdős–Rényi random graph $G(n,p_{n})$.
We now state our main result under the Chung-Lu model.
Theorem 3 (Semicircle law for the Chung-Lu model).
Consider the Chung-Lu random graph $G(n,\mathbf{w}_{n})$ defined above, and let
$w_{n,\min}:=\min_{1\leq i\leq n}w_{n,i},\qquad\bar{w}_{n}:=\frac{1}{n}\sum_{i=1}^{n}w_{n,i}.$
Assume that
$w_{n,\min}\gg\sqrt{\bar{w}_{n}}.$ (5)
Then the empirical spectral distribution of $\sqrt{\bar{w}_{n}}(\mathbf{I}_{n}-\mathcal{L}_{n})$ converges weakly in probability to the semicircle law.
Remark 2 (Comparison with [chung2003spectra]).
Condition (5) implies that $w_{n,\min}\to\infty$ as $n\to\infty$. Under the same assumptions, Chung, Lu, and Vu [chung2003spectra, Theorem 6] established a semicircle law only for the proxy matrix $\mathbf{W}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{W}_{n}^{-1/2}$. The present work closes this gap by establishing the semicircle law for the normalized Laplacian itself.
3. Preliminaries
Let $\mathbf{M}$ be a symmetric matrix with eigenvalues $\lambda_{1}(\mathbf{M})\geq\lambda_{2}(\mathbf{M})\geq\cdots\geq\lambda_{n}(\mathbf{M})$. The empirical spectral distribution (ESD) of $\mathbf{M}$, denoted by $\hat{\mu}(\mathbf{M})$, is defined as
$\hat{\mu}(\mathbf{M})=\frac{1}{n}\sum_{j=1}^{n}\delta_{\lambda_{j}(\mathbf{M})}.$
If $\mathbf{M}$ is random, then $\hat{\mu}(\mathbf{M})$ is a random probability measure on $(\mathbb{R},\mathcal{B})$. We write $\hat{\mu}(\mathbf{M}_{n})\Rightarrow\mu$ to denote that $\hat{\mu}(\mathbf{M}_{n})$ converges weakly to a probability measure $\mu$ on $(\mathbb{R},\mathcal{B})$.
Definition 1 (bounded Lipschitz metric).
Let $d_{\text{BL}}$ denote the bounded Lipschitz metric
$d_{\text{BL}}(\mu,\nu)=\sup\left\{\int f\,d\mu-\int f\,d\nu:\|f\|_{\infty}+\|f\|_{L}\leq 1\right\},$ (6)
where $\|f\|_{\infty}=\sup\limits_{x}|f(x)|$, $\|f\|_{L}=\sup\limits_{x\neq y}\dfrac{|f(x)-f(y)|}{|x-y|}$.
We will use the following comparison inequality to establish weak convergence of empirical spectral distributions:
Lemma 1.
For any symmetric matrices $A,B$, we have
$d_{\text{BL}}^{2}(\hat{\mu}(\mathbf{A}),\hat{\mu}(\mathbf{B}))\leq\frac{1}{n}\operatorname{tr}((\mathbf{B}-\mathbf{A})^{2}).$ (7)
Proof.
For the spectral measures of $n\times n$ symmetric real matrices $\mathbf{A},\mathbf{B}$ we have
$d_{\text{BL}}(\hat{\mu}(\mathbf{A}),\hat{\mu}(\mathbf{B}))$ $\leq\sup\left\{\frac{1}{n}\sum_{j=1}^{n}|f(\lambda_{j}(\mathbf{A}))-f(\lambda_{j}(\mathbf{B}))|:\|f\|_{L}\leq 1\right\}$
$\leq\frac{1}{n}\sum_{j=1}^{n}|\lambda_{j}(\mathbf{A})-\lambda_{j}(\mathbf{B})|$
$\leq\frac{1}{\sqrt{n}}\sqrt{\sum_{j=1}^{n}|\lambda_{j}(\mathbf{A})-\lambda_{j}(\mathbf{B})|^{2}}\leq\frac{1}{\sqrt{n}}\sqrt{\operatorname{tr}((\mathbf{B}-\mathbf{A})^{2})},$
where the last inequality is from Lidskii’s theorem (see, e.g., [bai1999methodologies, Lemma 2.3]). ∎
4. Proof of Theorems 1 and 2
Before proceeding, we briefly outline the overall proof strategy. The proof is divided into several steps. From Ding and Jiang [ding2010spectral], the empirical spectral distribution of $\mathbf{A}_{n}/\sqrt{np_{n}(1-p_{n})}$ converges weakly to the semicircle law when $np_{n}(1-p_{n})\to\infty$. Therefore, it suffices to show that the bounded Lipschitz distance between the empirical spectral distribution under consideration, and that of $\mathbf{A}_{n}/\sqrt{np_{n}(1-p_{n})}$ converges to zero almost surely as $n\to\infty$. A direct control of $d_{\text{BL}}$ is technically challenging. To address this issue, we decompose the analysis into two parts: the expectation term and the deviation from its expectation. The expectation is handled in Lemma 2, where we derive a suitable tail estimate, while the fluctuation around the expectation is controlled in Lemma 3 by an improved bounded differences inequality.
Lemma 2 is first used to establish Theorem 1. Then, by combining Lemmas 2 and 3, we show that the bounded Lipschitz distance converges to zero almost surely, which in turn yields the desired spectral convergence and completes the proof of Theorem 2.
Set
$\widetilde{\mathbf{\mathcal{L}}}_{n}=\mathbf{I}_{n}-\frac{1}{(n-1)p_{n}}\mathbf{A}_{n},\ u_{n}=(n-1)p_{n},\ d_{n,i}=\sum\limits_{j\neq i}a_{ij}^{(n)}.$ (8)
Note that
$\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}$ $=-\frac{1}{(n-1)p_{n}}\mathbf{A}_{n}+\operatorname{diag}(d_{n,1}^{-1/2},\ldots,d_{n,n}^{-1/2})\mathbf{A}_{n}\operatorname{diag}(d_{n,1}^{-1/2},\ldots,d_{n,n}^{-1/2})$
$=\left(\left(-\frac{1}{(n-1)p_{n}}+\frac{1}{\sqrt{d_{n,i}}\sqrt{d_{n,j}}}\right)a^{(n)}_{ij}\right)_{n\times n}.$
Lemma 2 (Approximation of $\frac{1}{n}\mathbb{E}(\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2})$).
Assume that the matrices $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$ are defined as (3) and (8). If $np_{n}\to\infty$, then
$\frac{1}{n}\mathbb{E}(\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2})=O(u_{n}^{-2}).$
Lemma 3 (Concentration).
Assume that the matrices $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$ are defined as (3) and (8) and $np_{n}\geq C\log n$. Then, asymptotically almost surely,
$\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}-\frac{1}{n}\mathbb{E}(\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2})=o(u_{n}^{-1}).$
4.1. Proof of Lemma 2
Through symmetry and the fact that $\{a_{ij}^{(n)}\}$ are i.i.d. random variables, we can derive that
$\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right]$ $=\frac{1}{n}\mathbb{E}\!\left[\sum_{i\neq j}^{n}\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,i}}\sqrt{d_{n,j}}}\right)^{2}a_{ij}^{(n)}\right]$ (9)
$=\frac{1}{n}\sum_{i\neq j}^{n}\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,i}}\sqrt{d_{n,j}}}\right)^{2}a_{ij}^{(n)}\right]$
$=\frac{1}{n}\cdot n(n-1)\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}a_{12}^{(n)}\right]$
$=(n-1)\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}a_{12}^{(n)}\right].$
Note that $\mathbb{P}(a_{12}^{(n)}=1)=1-\mathbb{P}(a_{12}^{(n)}=0)=p_{n}$, we get
$\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}a_{12}^{(n)}\right]=p_{n}\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}|a_{12}^{(n)}=1\right].$
We briefly outline our approach for evaluating
$\mathbb{E}\!\left[\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}d_{n,2}}}\right)^{2}\,\big|\,a_{12}^{(n)}=1\right].$
We introduce an event $E_{1}$ under which both $d_{n,1}$ and $d_{n,2}$ remain close to $u_{n}$, so that the function $(x,y)\mapsto\left(\frac{1}{u_{n}}-\frac{1}{\sqrt{xy}}\right)^{2}$ can be locally approximated. The expectation is then decomposed into two parts corresponding to $E_{1}$ and its complement $E_{1}^{c}$. For $E_{1}$, we use independence and moment estimates of binomial variables to show that the contribution is $O(u_{n}^{-3})$, while for $E_{1}^{c}$, a Chernoff bound ensures that its probability is exponentially small. Combining these estimates yields the desired order of magnitude for the entire expectation.
Given that $a_{12}^{(n)}=1$, let
$d_{n,1}=1+Z_{1}:=1+\sum_{k\neq 1,2}a_{1k}^{(n)},\qquad d_{n,2}=1+Z_{2}:=1+\sum_{k\neq 1,2}a_{2k}^{(n)}.$
So we have
$Z_{1}\sim\mathrm{Bin}(n-2,p_{n})\quad\text{and}\quad Z_{2}\sim\mathrm{Bin}(n-2,p_{n}),$
where $\mathrm{Bin}(n,p)$ denotes the binomial distribution with parameters $n$ and $p$. Thus
$\mathbb{E}(Z_{1}+1)=\mathbb{E}(Z_{2})=(n-2)p_{n}+1=u_{n}+o(u_{n}),$
$\mathbb{E}\left(\left(Z_{1}+1\right)^{2}\right)=\mathbb{E}\left(\left(Z_{2}+1\right)^{2}\right)=3(n-2)p_{n}+(n-2)(n-3)p_{n}^{2}+1=u_{n}^{2}+O(u_{n}).$
Let the event
$E_{1}:=\{d_{n,1},d_{n,2}\in[u_{n}/2,3u_{n}/2]\}.$
We first approximate
$\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}\mathbf{1}(E_{1})|a_{12}^{(n)}=1\right).$
Using the identity $|\frac{1}{\sqrt{x}}-\frac{1}{\sqrt{y}}|=|\frac{\sqrt{y}-\sqrt{x}}{\sqrt{xy}}|=|\frac{y-x}{\sqrt{xy}(\sqrt{x}+\sqrt{y})}|\leq|\frac{y-x}{y\sqrt{x}}|$ for $x=(Z_{1}+1)(Z_{2}+1)\ ,y=u_{n}^{2}$, we obtain
$\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}\mathbf{1}(E_{1})|a_{12}^{(n)}=1\right)=$ $\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{Z_{1}+1}\sqrt{Z_{2}+1}}\right)^{2}\mathbf{1}(E_{1})\right)$
$\leq$ $\mathbb{E}\!\left(\left(\frac{(Z_{1}+1)(Z_{2}+1)-u_{n}^{2}}{u_{n}^{2}\sqrt{Z_{1}+1}\sqrt{Z_{2}+1}}\right)^{2}\mathbf{1}(E_{1})\right).$
By non-negativity, note that $((Z_{1}+1)(Z_{2}+1)-u_{n}^{2})^{2}\mathbf{1}(E_{1})\leq((Z_{1}+1)(Z_{2}+1)-u_{n}^{2})^{2}$,and when event $E_{1}$ holds, $(Z_{i}+1)^{-1}\leq 2/u_{n}$ for $i=1,2$. Since $Z_{1}$ and $Z_{2}$ are independent, we can therefore bound the expectation by
$\mathbb{E}\!\left(\left(\frac{(Z_{1}+1)(Z_{2}+1)-u_{n}^{2}}{u_{n}^{2}\sqrt{Z_{1}+1}\sqrt{Z_{2}+1}}\right)^{2}\mathbf{1}(E_{1})\right)\leq$ $\mathbb{E}\!\left(\frac{4((Z_{1}+1)(Z_{2}+1)-u_{n}^{2})^{2}}{u_{n}^{6}}\right)$ (10)
$=$ $\frac{4}{u_{n}^{6}}\cdot\mathbb{E}\!\left(\left(Z_{1}+1\right)^{2}\left(Z_{2}+1\right)^{2}-2u_{n}^{2}(Z_{1}+1)(Z_{2}+1)+u_{n}^{4}\right)$
$=$ $\frac{4}{u_{n}^{6}}\cdot\left(\mathbb{E}\!\left(\left(Z_{1}+1\right)^{2}\right)\mathbb{E}\!\left(\left(Z_{2}+1\right)^{2}\right)-2u_{n}^{2}\mathbb{E}\!\left(Z_{1}+1\right)\mathbb{E}\!\left(Z_{2}+1\right)+u_{n}^{4}\right)$
$=$ $\frac{4}{u_{n}^{6}}(u_{n}^{4}+O(u_{n}^{3})-2u_{n}^{4}+o(u_{n}^{3})+u_{n}^{4})$
$=$ $O(u_{n}^{-3}).$
Under the assumption $np_{n}\to\infty$, a direct application of Chernoff’s bound yields
$\mathbb{P}(E_{1}^{c}\mid a_{12}=1)\leq 2\exp\!\left(-u_{n}/12\right).$ (11)
When event $E_{1}^{c}$ occurs, observe that
$|\frac{1}{u_{n}}-\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}|\leq|\frac{1}{u_{n}}|+|\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}|\leq 1+1=2,$
where we use the pseudo-inverse convention of $\mathbf{D}_{n}^{-1}$.
Thus we have
$\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}\mathbf{1}(E_{1}^{C})|a_{12}^{(n)}=1\right)\leq 4\mathbb{P}(E_{1}^{c})\leq 8\exp\!\left(-u_{n}/12\right).$ (12)
Since $u_{n}\to\infty$ as $n\to\infty$, it follows that
$\exp\!\left(-u_{n}/12\right)=o\!\left(u_{n}^{-3}\right).$
Combine (10) and (12), we get
$\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}|a_{12}^{(n)}=1\right)=$ $\mathbb{E}\!\left(\left(-\frac{1}{u_{n}}+\frac{1}{\sqrt{d_{n,1}}\sqrt{d_{n,2}}}\right)^{2}\left(\mathbf{1}(E_{1})+\mathbf{1}(E_{1}^{c})\right)|a_{12}^{(n)}=1\right)$ (13)
$=$ $O(u_{n}^{-3}).$
Substituting (13) into (9), we obtain
$\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right]=(n-1)p_{n}\cdot O(u_{n}^{-3})=O(u_{n}^{-2}).$ (14)
4.2. Proof of Theorem 1
We begin by comparing the empirical spectral distributions (ESDs) of $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$. Since the asymptotic spectral behavior of $\widetilde{\mathbf{\mathcal{L}}}_{n}$ has already been established in Ding and Jiang [ding2010spectral], it is enough to show that
$d_{\mathrm{BL}}\!\left(\hat{\mu}(\mathbf{\mathcal{L}}_{n}),\hat{\mu}(\widetilde{\mathbf{\mathcal{L}}}_{n})\right)\longrightarrow 0\quad\text{in probability as }n\to\infty.$
By Lemma 2, we have
$\mathbb{E}\!\left[d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\widetilde{\mathbf{\mathcal{L}}}_{n}),\hat{\mu}(\mathbf{\mathcal{L}}_{n})\right)\right]\leq\frac{1}{n}\,\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right]=O(u_{n}^{-2}).$ (15)
By Markov’s inequality, (15) implies that the ESDs of $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$ are asymptotically indistinguishable in the $d_{\mathrm{BL}}$ metric. The same argument applies to scaled and shifted versions of $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$. Indeed, for any sequences of constants $\{\gamma_{n}\}$ and $\{\rho_{n}\}$,
$\mathbb{E}\!\left[d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n}),\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\right)\right]=O\!\left(\frac{\gamma_{n}^{2}}{u_{n}^{2}}\right).$ (16)
Applying Markov’s inequality yields
$\mathbb{P}\!\left(d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n}),\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\right)\geq\varepsilon\right)\leq c\,\frac{\gamma_{n}^{2}}{u_{n}^{2}\,\varepsilon}.$
Now choose
$\gamma_{n}=-\rho_{n}=-\sqrt{np_{n}/(1-p_{n})}.$ (17)
Then
$\frac{\gamma_{n}^{2}}{u_{n}^{2}}=O\!\left(\frac{1}{np_{n}(1-p_{n})}\right)=o(1),$
and hence
$\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n})\Longrightarrow\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\quad\text{in probability as }n\to\infty.$ (18)
By the result of Ding and Jiang [ding2010spectral], we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\!\left(\sqrt{\frac{np_{n}}{1-p_{n}}}\left(1-\lambda_{i}(\widetilde{\mathbf{\mathcal{L}}}_{n})\right)\leq x\right)\Longrightarrow\mu_{\mathrm{sc}},\quad\text{a.s. as }n\to\infty.$
With (17), for all $1\leq i\leq n$,
$\lambda_{n-i}\!\left(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n}\right)=\sqrt{\frac{np_{n}}{1-p_{n}}}\left(1-\lambda_{i}(\widetilde{\mathbf{\mathcal{L}}}_{n})\right).$
Therefore,
$\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\Longrightarrow\mu_{\mathrm{sc}}\quad\text{a.s. as }n\to\infty.$ (19)
Finally, combining (18) and (19) yields
$\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n})\Longrightarrow\mu_{\mathrm{sc}}\quad\text{in probability as }n\to\infty,$
which completes the proof.
4.3. Proof of Lemma 3
Firstly, we introduce the following result of Warnke [warnke2016method].
Lemma 4 (Typical bounded differences inequality).
Let $X=(X_{1},\ldots,X_{N})$ be a family of independent random variables with $X_{k}$ taking values in a set $\Lambda_{k}$. Let $\Gamma\subseteq\prod_{j\in[N]}\Lambda_{j}$ be an event, and assume that the function $F:\prod_{j\in[N]}\Lambda_{j}\to\mathbb{R}$ satisfies the following typical Lipschitz condition:
(TL) There exist sequences $(c_{k})_{k\in[N]}$ and $(d_{k})_{k\in[N]}$ with $c_{k}\leq c_{k}^{\prime}$ such that whenever $x,\sigma_{k}(x)\in\prod_{j\in[N]}\Lambda_{j}$ differ only in the $k$-th coordinate, we have
$|F(x)-F(\sigma_{k}(x))|\leq\begin{cases}c_{k},&\text{if }x\in\Gamma,\\[3.0pt] c_{k}^{\prime},&\text{otherwise.}\end{cases}$
For any $(\beta_{k})_{k\in[N]}$ with $\beta_{k}\in(0,1]$, there exists an event $\mathcal{A}=\mathcal{A}(\Gamma,(\beta_{k})_{k\in[N]})$ satisfying
$\mathbb{P}(\mathcal{A})\leq\sum_{k\in[N]}\beta_{k}^{-1}\,\mathbb{P}(X\notin\Gamma),\qquad\neg\mathcal{A}\subseteq\Gamma,$ (20)
such that, for $\mu=\mathbb{E}F(X)$, $\varepsilon_{k}=\beta_{k}(c_{k}^{\prime}-c_{k})$, and any $t\geq 0$,
$\mathbb{P}(F(X)\geq\mu+t\text{ and }\neg\mathcal{A})\leq\exp\!\left(-\frac{t^{2}}{2\sum_{k\in[N]}(c_{k}+\varepsilon_{k})^{2}}\right).$
Proof of Lemma 3.
Consider the function $f_{n}:\{0,1\}^{n\times n}\to\mathbb{R}$ defined by
$f_{n}(\{X_{i,j}\}_{1\leq i,j\leq n}):=\sum_{i\neq j}^{n}\left(u_{n}^{-1}-\left(\sum_{k\neq i}X_{ik}\sum_{k\neq j}X_{jk}\right)^{-1/2}\right)^{2}X_{ij},$ (21)
which depends only on the adjacency variables $\{X_{ij}\}$ of the random graph. Then
$F=\frac{1}{n}f_{n}=\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}.$
If we change only one coordinate $(u,v)$ of $\{X_{ij}\}$ while keeping all others fixed, define
$\sigma_{uv}(\{X_{ij}\}):=\begin{cases}1-X_{ij},&\text{if }(i,j)=(u,v)\text{ or }(i,j)=(v,u),\\[3.0pt] X_{ij},&\text{otherwise.}\end{cases}$
Let $S_{l}=\sum_{k\neq l}^{n}X_{kl}$ and $S_{l}^{\prime}=\sum_{k\neq l}^{n}\sigma_{uv}(\{X_{ij}\})_{kl}$. It is easy to verify that
$|S_{k}-S_{k}^{\prime}|=\begin{cases}1,&\text{if }k=u\text{ or }k=v,\\[3.0pt] 0,&\text{otherwise.}\end{cases}$
By definition of $f_{n}$ in (21), we obtain
$\delta_{uv}(\{X_{i,j}\})$ $=f_{n}(\{X_{i,j}\})-f_{n}(\sigma_{uv}(\{X_{ij}\}))$
$=2\sum_{k\neq u,v}^{n}\!\left[\!\left(u_{n}^{-1}-(S_{u}S_{k})^{-1/2}\right)^{2}-\left(u_{n}^{-1}-(S_{u}^{\prime}S_{k})^{-1/2}\right)^{2}\!\right]\!X_{uk}$
$\quad+2\sum_{k\neq u,v}^{n}\!\left[\!\left(u_{n}^{-1}-(S_{v}S_{k})^{-1/2}\right)^{2}-\left(u_{n}^{-1}-(S_{v}^{\prime}S_{k})^{-1/2}\right)^{2}\!\right]\!X_{vk}$
$\quad+2\!\left[\left(u_{n}^{-1}-(S_{u}S_{v})^{-1/2}\right)^{2}-\left(u_{n}^{-1}-(S_{u}^{\prime}S_{v}^{\prime})^{-1/2}\right)^{2}\right].$
Define the “typical” event
$\Gamma:=\left\{d_{n,i}\in\left(\frac{u_{n}}{2},\frac{3u_{n}}{2}\right)\text{ for all }1\leq i\leq n\right\}.$
We now estimate the typical Lipschitz constants $c_{k}$.
Consider the auxiliary function $g(x,y)=\left(u_{n}^{-1}-\frac{1}{\sqrt{xy}}\right)^{2}$. A direct computation yields
$\frac{\partial g}{\partial x}=2\!\left(u_{n}^{-1}-\frac{1}{\sqrt{xy}}\right)\!\left(-\frac{1}{2}x^{-3/2}y^{-1/2}\right)=\left(u_{n}^{-1}-\frac{1}{\sqrt{xy}}\right)x^{-3/2}y^{-1/2}.$
For $x,y\in(u_{n}/2,3u_{n}/2)$, we have $x^{-3/2}y^{-1/2}\leq 4u_{n}^{-2}$ and $\left|u_{n}^{-1}-x^{-1/2}y^{-1/2}\right|\leq u_{n}^{-1}$, hence $|\partial g/\partial x|\leq 4u_{n}^{-3}$. By symmetry, the same holds for $\partial g/\partial y$. Consequently, if both $x$ and $y$ change by at most $1$, the function $g(x,y)$ changes by at most $cu_{n}^{-3}$.
When the event $\Gamma$ holds, changing one coordinate $(u,v)$ in $\{X_{ij}\}$ affects at most the degrees of vertices $u$ and $v$. Recalling that $d_{n,i}=\sum_{j\neq i}a_{ij}$, we have
$|\delta_{uv}(\mathbf{A}_{n})|$ $\leq 2\sum_{k\neq u,v}^{n}cu_{n}^{-3}a_{uk}+2\sum_{k\neq u,v}^{n}cu_{n}^{-3}a_{vk}+cu_{n}^{-3}$
$\leq cu_{n}^{-3}(d_{n,u}+d_{n,v}+1)\leq 3cu_{n}^{-2},$
since $d_{n,i}\in[u_{n}/2,3u_{n}/2]$ under $\Gamma$. Hence, for $F:=n^{-1}f_{n}$, the typical Lipschitz constants satisfy
$c_{uv}\leq\frac{C_{1}}{n\,u_{n}^{2}},$
whenever the event $\Gamma$ holds.
Outside the typical event $\Gamma$, we distinguish two cases.
First, if $S_{u}=0$ or $S_{u}^{\prime}=0$, then necessarily $X_{uk}=0$ for all $k\neq u,v$. In this case the corresponding contribution vanishes identically, namely,
$\left|\sum_{k\neq u,v}^{n}\Big[\big(u_{n}^{-1}-(S_{u}S_{k})^{-1/2}\big)^{2}-\big(u_{n}^{-1}-(S_{u}^{\prime}S_{k})^{-1/2}\big)^{2}\Big]X_{uk}\right|=0.$
Otherwise, assume that $S_{u}S_{u}^{\prime}\neq 0$. Using the identities $x^{2}-y^{2}=(x+y)(x-y)$ and
$x^{-1/2}-y^{-1/2}=\frac{x-y}{xy\big(x^{-1/2}+y^{-1/2}\big)},$
we obtain
$\left|\sum_{k\neq u,v}^{n}\Big[\big(u_{n}^{-1}-(S_{u}S_{k})^{-1/2}\big)^{2}-\big(u_{n}^{-1}-(S_{u}^{\prime}S_{k})^{-1/2}\big)^{2}\Big]X_{uk}\right|$
$\leq$ $\left|\sum_{k\neq u,v}^{n}\Big(2u_{n}^{-1}-(S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)\Big((S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)X_{uk}\right|$
$\leq$ $c\left|\sum_{k\neq u,v}^{n}\Big((S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)X_{uk}\right|$
$\leq$ $c\,\max_{k}S_{k}^{-1/2}\frac{|S_{u}-S_{u}^{\prime}|}{S_{u}S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}\sum_{k\neq u,v}^{n}X_{uk}$
$\leq$ $c\,\max_{k}S_{k}^{-1/2}\frac{S_{u}}{S_{u}S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}$
$=$ $c\,\max_{k}S_{k}^{-1/2}\frac{1}{S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}.$
Under the pseudo-inverse convention, if there exists $1\leq i\leq n$ such that $d_{n,i}=O(1)$, then the increment $\delta_{uv}$ may attain its maximal size, and we obtain the crude bound
$\delta_{uv}=O(1).$
Therefore, recall that $F:=n^{-1}f_{n}$, we also have the crude worst-case bound
$c^{\prime}_{uv}\leq\frac{C_{2}}{n},$
for some absolute constant $C_{2}>0$.
We choose the parameters uniformly over edges $k\leftrightarrow(u,v)$ by
$\beta_{k}=\beta_{n}:=u_{n}^{-2}\in(0,1]\quad\text{for all sufficiently large }n,$
so that $\varepsilon_{k}=\beta_{n}(c_{k}^{\prime}-c_{k})\leq C_{2}n^{-1}u_{n}^{-2}$. Since $u_{n}\to\infty$ under the assumption $np_{n}\gtrsim\log n$, the typical constants dominate:
$c_{k}+\varepsilon_{k}\leq(C_{1}+C_{2})n^{-1}u_{n}^{-2}.$
Summing over the $N=\binom{n}{2}$ independent edges gives
$\sum_{k}(c_{k}+\varepsilon_{k})^{2}\leq N\cdot\frac{(C_{1}+C_{2})^{2}}{n^{2}u_{n}^{4}}=O(u_{n}^{-4}).$
Therefore, by (20), we obtain the sub-Gaussian tail bound
$\mathbb{P}(F-\mathbb{E}F\geq t\text{ and }\neg\mathcal{A})\leq\exp(-c\,t^{2}u_{n}^{4}).$ (22)
for some absolute constant $c>0$.
It remains to control $\mathbb{P}(\mathcal{A})$. Since $d_{n,i}\sim\mathrm{Bin}(n-1,p_{n})$ and $u_{n}=(n-1)p_{n}$, Chernoff’s bound gives
$\mathbb{P}(|d_{n,i}-u_{n}|\geq u_{n}/2)\leq 2e^{-u_{n}/12}.$
By a union bound,
$\mathbb{P}(\neg\Gamma)\leq 2ne^{-u_{n}/12}.$
This also implies $\mathbf{D}_{n}$ is invertible asymptotically almost surely. Using $\beta_{n}=u_{n}^{-2}$, we deduce from (20) that
$\mathbb{P}(\mathcal{A})\leq N\beta_{n}^{-1}\mathbb{P}(\neg\Gamma)\lesssim n^{3}u_{n}^{2}e^{-u_{n}/12},$ (23)
which is summable in $n$ under $u_{n}\gtrsim\log n$ (equivalently, $np_{n}\gtrsim\log n$).
Finally, set $t_{n}=k/u_{n}$ with any fixed $k>0$. By (22) and (23),
[…]
By the Borel–Cantelli lemma,
$\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}-\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}\right]=o\!\left(\frac{1}{u_{n}}\right)\qquad\text{a.s. as }n\to\infty,$
which establishes the desired concentration result. ∎
4.4. Proof of Theorem 2
Proof.
It suffices to show that
$d_{\mathrm{BL}}\!\left(\hat{\mu}(\mathbf{\mathcal{L}}_{n}),\hat{\mu}(\widetilde{\mathbf{\mathcal{L}}}_{n})\right)\longrightarrow 0\quad\text{a.s. as }n\to\infty.$
By Lemmas 2 and 3, we have
$d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\widetilde{\mathbf{\mathcal{L}}}_{n}),\hat{\mu}(\mathbf{\mathcal{L}}_{n})\right)$ $\leq\frac{1}{n}\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}$
$\leq\Biggl|\frac{1}{n}\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}-\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right]\Biggr|+\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right].$
Lemma 2 yields
$\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\!\left(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n}\right)^{2}\right]=O(u_{n}^{-2}),$
while Lemma 3 implies that the centered fluctuation term is $o(u_{n}^{-1})$ almost surely. Consequently,
$d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\widetilde{\mathbf{\mathcal{L}}}_{n}),\hat{\mu}(\mathbf{\mathcal{L}}_{n})\right)=o(u_{n}^{-1})\quad\text{a.s. as }n\to\infty.$ (24)
(24) is also true under a linear scaling of $\{\gamma_{n}\}$ and $\{\rho_{n}\}$, we have
$d_{\mathrm{BL}}^{2}\!\left(\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n}),\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\right)=o\!\left(\frac{\gamma_{n}^{2}}{u_{n}}\right)\quad\text{a.s. as }n\to\infty.$ (25)
On the other hand, recall (19), we know that
$\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\Longrightarrow\mu_{\mathrm{sc}}\quad\text{a.s. as }n\to\infty,$
which is equivalent to
$d_{\mathrm{BL}}\!\left(\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n}),\mu_{\mathrm{sc}}\right)\longrightarrow 0\quad\text{a.s. as }n\to\infty.$
Note that
$\gamma_{n}=-\rho_{n}=-\sqrt{\frac{np_{n}}{1-p_{n}}}.$
Under the assumption $np_{n}\geq C\log n$, we have $u_{n}\asymp np_{n}$ and hence
$\frac{\gamma_{n}^{2}}{u_{n}}=O(1).$
Therefore, by the triangle inequality for $d_{\mathrm{BL}}$,
$d_{\mathrm{BL}}\!\left(\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n}),\mu_{\mathrm{sc}}\right)$ $\leq d_{\mathrm{BL}}\!\left(\hat{\mu}(\gamma_{n}\mathbf{\mathcal{L}}_{n}+\rho_{n}\mathbf{I}_{n}),\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n})\right)$
$\quad+d_{\mathrm{BL}}\!\left(\hat{\mu}(\gamma_{n}\widetilde{\mathbf{\mathcal{L}}}_{n}+\rho_{n}\mathbf{I}_{n}),\mu_{\mathrm{sc}}\right)\longrightarrow 0\quad\text{a.s. as }n\to\infty.$
This completes the proof of Theorem 2. ∎
5. Proof of Theorem 3
Proof.
Define the normalized adjacency matrices
$\mathbf{N}_{n}:=\mathbf{D}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{D}_{n}^{-1/2},\qquad\mathbf{\widetilde{N}}_{n}:=\mathbf{W}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{W}_{n}^{-1/2}.$
We proceed in two steps. First, we establish the semicircle law for the matrix $\mathbf{\widetilde{N}}_{n}$. Second, we show that the same limiting law holds for $N_{n}$ by a perturbation analysis.
Recall that Chung, Lu, and Vu [chung2003spectra] proved that, under condition (5), the normalized matrix
$\mathbf{C}_{\mathrm{nor}}:=\sqrt{\bar{w}_{n}}\,\mathbf{C},\qquad\mathbf{C}:=\mathbf{W}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{W}_{n}^{-1/2}-\phi_{n}W_{n}^{1/2}\mathbf{KW}_{n}^{1/2},$
has an empirical spectral distribution converging in probability to the semicircle distribution $\mu_{sc}$.
Observe that
$\mathbf{\widetilde{N}_{n}}=\mathbf{W}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{W}_{n}^{-1/2}=\mathbf{C}+\phi_{n}\mathbf{W}_{n}^{1/2}\mathbf{KW}_{n}^{1/2}=:\mathbf{C}+\mathbf{R}_{W},$
where $\mathbf{K}$ denotes the all-ones matrix. The matrix $\mathbf{R}_{W}$ has rank one, since
$\mathbf{W}_{n}^{1/2}\mathbf{KW}_{n}^{1/2}=\bigl(\sqrt{w_{n,1}},\dots,\sqrt{w_{n,n}}\bigr)^{\!\top}\bigl(\sqrt{w_{n,1}},\dots,\sqrt{w_{n,n}}\bigr).$
Consequently, $\mathbf{\widetilde{N}}_{n}$ differs from $\mathbf{C}$ by a rank-one perturbation.
We now invoke the rank inequality due to Bai [bai2010spectral, Lemma 2.6].
Lemma 5 (Rank inequality).
Let $\mathbf{A}$ and $\mathbf{B}$ be two $n\times n$ Hermitian matrices. Then
$\bigl\|\hat{\mu}(\mathbf{A})-\hat{\mu}(\mathbf{B})\bigr\|_{\infty}\leq\frac{1}{n}\,\operatorname{rank}(\mathbf{A-B}).$
Applying Lemma 5 yields
$\bigl\|\hat{\mu}(\mathbf{\widetilde{N}}_{n})-\hat{\mu}(\mathbf{C})\bigr\|_{\infty}\leq\frac{1}{n},$ (26)
which implies that $\mathbf{\widetilde{N}}_{n}$ and $\mathbf{C}$ have the same limiting spectral distribution.
Combining (26) observation with the semicircle law for $\mathbf{C}_{\mathrm{nor}}$, we conclude that
$\hat{\mu}\!\left(\sqrt{\bar{w}_{n}}\,\mathbf{\widetilde{N}}_{n}\right)\;\Longrightarrow\;\mu_{\mathrm{sc}}\quad\text{in probability as }n\to\infty.$ (27)
Lemma 6.
There exist absolute constants $c,c^{\prime}>0$ such that
$\frac{1}{n}\,\mathbb{E}\,\operatorname{tr}\!\left((N_{n}-\widetilde{N}_{n})^{2}\right)\;\leq\;\frac{c}{\bar{w}_{n}\,w_{n,\min}}\;+\;c\,\bar{w}_{n}\,e^{-c^{\prime}w_{n,\min}}.$
In particular, since $w_{n,\min}\to\infty$, we have
$\bar{w}_{n}\cdot\frac{1}{n}\,\mathbb{E}\,\operatorname{tr}\!\left((\mathbf{N}_{n}-\mathbf{\widetilde{N}}_{n})^{2}\right)=O\!\left(\frac{1}{w_{n,\min}}\right)\to 0\quad\text{in probability as }n\to\infty.$
Proof.
For $i\neq j$, the entries of $\mathbf{N}_{n}$ and $\mathbf{\widetilde{N}}_{n}$ are given by
$(\mathbf{N}_{n})_{ij}=\frac{a_{ij}^{(n)}}{\sqrt{d_{n,i}d_{n,j}}},\qquad(\mathbf{\widetilde{N}}_{n})_{ij}=\frac{a_{ij}^{(n)}}{\sqrt{w_{n,i}w_{n,j}}}.$
Hence,
$\mathbf{\widetilde{N}}_{n}-\mathbf{N}_{n}=\Bigl(\Bigl(\frac{1}{\sqrt{w_{n,i}w_{n,j}}}-\frac{1}{\sqrt{d_{n,i}d_{n,j}}}\Bigr)a_{ij}^{(n)}\Bigr)_{1\leq i,j\leq n}.$
By symmetry and independence of the edge indicators,
$\frac{1}{n}\mathbb{E}\operatorname{tr}\!\left((\mathbf{\widetilde{N}}_{n}-\mathbf{N}_{n})^{2}\right)$ $=\frac{1}{n}\sum_{i\neq j}\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{w_{n,i}w_{n,j}}}-\frac{1}{\sqrt{d_{n,i}d_{n,j}}}\Bigr)^{2}a_{ij}^{(n)}\right].$ (28)
Fix $i\neq j$. By conditioning on $a_{ij}^{(n)}$,
$\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}a_{ij}^{(n)}\right]=p_{n,ij}\,\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}\,\Big|\,a_{ij}^{(n)}=1\right].$
Define the event
$E_{3}:=\Bigl\{d_{n,i}\in[\tfrac{1}{2}w_{n,i},\,\tfrac{3}{2}w_{n,i}],\;d_{n,j}\in[\tfrac{1}{2}w_{n,j},\,\tfrac{3}{2}w_{n,j}]\Bigr\}.$
On $E_{3}$, a Taylor expansion argument yields
$\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}\mathbf{1}_{E_{3}}\,\Big|\,a_{ij}^{(n)}=1\right]\leq c\!\left(\frac{1}{w_{n,i}^{2}w_{n,j}}+\frac{1}{w_{n,j}^{2}w_{n,i}}\right).$
Since $d_{n,i}$ and $d_{n,j}$ are sums of independent Bernoulli random variables, by a Chernoff bound, we have
$\mathbb{P}(E_{3}^{c})\leq 2e^{-w_{n,i}/12}+2e^{-w_{n,j}/12}\leq 4e^{-w_{n,\min}/12}.$
On $E_{3}^{c}$, we use the crude bound and the pseudo-inverse convention to have
$\Bigl|\frac{1}{\sqrt{w_{n,i}w_{n,j}}}-\frac{1}{\sqrt{d_{n,i}d_{n,j}}}\Bigr|\leq\frac{1}{\sqrt{w_{n,i}w_{n,j}}}+\frac{1}{\sqrt{d_{n,i}d_{n,j}}}\leq 2,$
which implies
$\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}\mathbf{1}_{E_{3}^{c}}\,\Big|\,a_{ij}^{(n)}=1\right]\leq 16e^{-w_{n,\min}/12}.$
Combining the two cases, we obtain
$\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}\,\Big|\,a_{ij}^{(n)}=1\right]\leq c\!\left(\frac{1}{w_{n,i}^{2}w_{n,j}}+\frac{1}{w_{n,j}^{2}w_{n,i}}\right)+16e^{-w_{n,\min}/12}.$
Multiplying by $p_{n,ij}=\phi_{n}w_{n,i}w_{n,j}$ gives
$\mathbb{E}\!\left[\Bigl(\frac{1}{\sqrt{d_{n,i}d_{n,j}}}-\frac{1}{\sqrt{w_{n,i}w_{n,j}}}\Bigr)^{2}a_{ij}^{(n)}\right]\leq c\phi_{n}\!\left(\frac{1}{w_{n,i}}+\frac{1}{w_{n,j}}\right)+16p_{n,ij}e^{-w_{n,\min}/12}.$
Summing over $i\neq j$ and using $\sum_{i,j}p_{n,ij}=n\bar{w}_{n}=1/\phi_{n}$, we conclude from (28) that
$\frac{1}{n}\mathbb{E}\operatorname{tr}\!\left((\mathbf{\widetilde{N}}_{n}-\mathbf{N}_{n})^{2}\right)\leq\frac{c}{\bar{w}_{n}w_{n,\min}}+c\bar{w}_{n}e^{-w_{n,\min}/12},$
which completes the proof. ∎
Recall that, for any sequences $\gamma_{n}\in\mathbb{R}$, we have
$d_{BL}^{2}\left(\hat{\mu}(\gamma_{n}\mathbf{N}_{n}),\hat{\mu}(\gamma_{n}\mathbf{\widetilde{N}}_{n})\right)\leq\gamma_{n}^{2}\cdot\frac{1}{n}\operatorname{tr}\left((\mathbf{N}_{n}-\mathbf{\widetilde{N}}_{n})^{2}\right),$
Taking $\gamma_{n}=\sqrt{\bar{w}_{n}}$, and using the expectation value and Lemma 6, we get:
$\mathbb{E}\left[d_{BL}^{2}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n})\right)\right]\leq\bar{w}_{n}\cdot\frac{1}{n}\mathbb{E}\operatorname{tr}\left((\mathbf{N}_{n}-\mathbf{\widetilde{N}}_{n})^{2}\right)=O\left(\frac{1}{w_{n,\min}}\right)\to 0\quad\text{as n}\to\infty,$
where we use the assumption $w_{n,\min}\gg\sqrt{\overline{w}_{n}}$. Thus, by Markov’s inequality, we obtain
$d_{BL}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n})\right)\to 0\quad\text{in probability as }n\to\infty.$
Thus, by (27) and the triangle inequality,
$d_{BL}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\mu_{\text{sc}}\right)$ $\leq d_{BL}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n})\right)$ (29)
$\quad+d_{BL}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n}),\mu_{\text{sc}}\right)$
$\longrightarrow 0\quad\text{in probability as }n\to\infty.$
And (29) proves
$\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{D}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{D}_{n}^{-1/2})\Rightarrow\mu_{\text{sc}}\quad\text{in probability as }n\to\infty$
as desired.
Acknowledgments
Y.Z. was partially supported by the Simons Grant MPS-TSM-00013944.